OpenStudy (anonymous):

Convert ellipse to its standard equation. x^2-6x+3y^2-6y=0 (the standard equation is (x-h)^2/a^2 - (y-k)^2/b^2 = 1

OpenStudy (anonymous):

x^2-6x+9-9+3(y^2-2y+1-1)=0 (x-3)^2-9+3((y-1)^2-1)=0 (x-3)^2-9+3(y-1)^2-3=0 (x-3)^2+3(y-1)^2=12 (x-3)^2/12+3(y-1)^2/12=1 (x-3)^2/12+(y-1)^2/4=1

OpenStudy (anonymous):

what I got was: (x^2-6x) + (3y^2-6y) = 0 (x-3)^2 - 3(y^2+2y) = 0 (x-3)^2 - 3(y+1)^2 how did you get your answer?

OpenStudy (tkhunny):

Brush up on Completing the Square. You can't just throw numbers in there. You must fix it after you break it. (x^2-6x) + (3y^2-6y) = 0 (x^2-6x + -----) + (3y^2-6y) = 0 + _____ (x^2-6x + 9) + (3y^2-6y) = 0 + 9 (x-3)^2 + (3y^2-6y) = 9 (x-3)^2 + 3(y^2-2y + _____) = 9 + 3(_____) (x-3)^2 + 3(y^2-2y + 1) = 9 + 3(1) (x-3)^2 + 3(y-1)^2 = 9 + 3 (x-3)^2 + 3(y-1)^2 = 12 \(\dfrac{(x-3)^{2}}{12} + \dfrac{(y-1)^{2}}{4} = 1\)